Grade 12th PassDiscuss with colleagues and IITians

If a,b,c are positive integers, then find the number of solutions of the equation

∑a + ∑ab = 4198 - abc, where ∑ denotes summation of terms, one at a time and two at a time respectively.

Pls provide solution with explanation

Speed Racer
14 Years agoGrade 12th Pass

2 Answers

ARITRA  SINHA
ApprovedApproved Tutor Answer14 Years ago

these are good questions which u should not ask for the solution of. try urself no matter how much time it takes.

Ashwin Muralidharan IIT Madras
ApprovedApproved Tutor Answer14 Years ago

Hi Speed Racer.

 

See here that,

a+b+c+ab+bc+ca = 4198 - abc

 

Now, (a+1)(b+1)(c+1) = a+b+c+ab+bc+ca+abc+1.

 

So the above equation can be written as (a+1)(b+1)(c+1) = 4199.

 

Now factorise 4199... 4199 = 13*17*19.

 

So the triplet is (12,16,18)

 

Which can be arranged in 3! ways for a,b,c.

 

Hence 6 Solutions.

 

Best Regards,

Ashwin (IIT Madras).